• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Ibrahim

    Log entry time 00:01:30 on November 6,2008

    Entry number 247031

    keyword=Swing Shift Summary

    Swing Shift Summary

    16:00 Shift started taking data on the reference cell (H2 at 135 psig).

    16:30 The RC took over to do a beam position check

    17:40 Back to the reference cell (H2 at 135 psig).

    18:20 The RC asked for controlled access.

    19:05 Back from controlled access

    19:05 Moved the target to empty and asked for tune beam. The MCC need 15min to do this.

    19:25 Moved the target to the optics position

    19:38 MCC are giving hall B tune beam and they cannot deliver CW beam to hall A for now.

    19:50 Beam is back (CW, 5uA, raster off)

    20:00 Noticed that BPMB is off. Called the MCC to set it.


    20:15 MCC still having difficulty setting the beam position within 0.2mm of what is requested,

    20:35 Asked MCC to do a harp scan.

    21:30 Moved to the Reference Cell (H2 at 135 psig). Will stay here for the next 2 hours (30 min runs).

    00:00 Shift ends while taking coincedence runs on H2. An additional 30min run should be taken by the owl shift.